OpenStudy (anonymous):

Integrate 1/(xlogx) dx.

OpenStudy (anonymous):

put t=ln x

hartnn (hartnn):

then dt would be dt=(1/x)dx and your integral will reduce to integral of 1/t dt ok?

OpenStudy (lgbasallote):

integration by parts is fun too...

OpenStudy (lgbasallote):

complicated..but fun

OpenStudy (anonymous):

\[\int\limits\limits\limits\limits_{}^{}\frac{ 1 }{ t }dt = \int\limits\limits\limits\limits_{}^{}\frac{ 1 }{ xlnx } dx \neq \int\limits_{}^{} \frac{ 1 }{ xlogx }\] What am I missing?

OpenStudy (anonymous):

in ur question it's ln x or log x?

OpenStudy (anonymous):

logx

OpenStudy (anonymous):

then change into natural log 1st ln x = 2.303 log x

OpenStudy (anonymous):

then do the substitution t=lnx

OpenStudy (anonymous):

does it make sense?

OpenStudy (anonymous):

or log x= log e * ln x

OpenStudy (anonymous):

so u can substitute t = log x also then dt = log e /x dx

OpenStudy (anonymous):

I'm getting \[\log x = \frac{ 1 }{ \ln10} lnx\]

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Ok here's what I have, would you mind just double checking it for me? \[\int\limits_{}^{}\frac{ 1 }{ xlogx } dx\] \[t=logx, dt=\frac{ 1 }{ xln10}\]\[\ln10\int\limits_{}^{}1/t dt = \ln10\ln \left| u \right|+c= \ln10\ln \left| logx \right|+c\] look good?